# Rethinking Approximate Gaussian Inference in Classification

### v1

@@ Appendix I Benchmark Metrics
I.1 Log Probability Proper Scoring Rule for the Predictive
First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on (Mucsányi et al., 2023).
@@ I.1 Log Probability Proper Scoring Rule for the Predictive
I.2 Expected Calibration Error
To set up the required quantities for the Expected Calibration Error (ECE) metric (Naeini et al., 2015), we follow the steps below, based on (Mucsányi et al., 2023).
@@ I.2 Expected Calibration Error
The Expected Calibration Error (ECE) metric (Naeini et al., 2015) is then defined as
@@
$\mathrm{conf}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{n\in B_{m}}\max_{c\in\{1,\dotsc,C\}}f_{c}(x_{n}).$ (89)
@@
I.3 Binary Log Probability Proper Scoring Rule for Correctness Prediction
@@
$\ell\equiv\ell(x,y)=1\left(\max_{c\in\{1,\dotsc,C\}}h_{c}(x)=y\right).$ (90)
@@
One can show that this is indeed a strictly proper scoring rule (Mucsányi et al., 2023).
I.4 Accuracy
@@
I.5 Area Under the Receiver Operating Characteristic Curve for Out-of-Distribution Detection
@@
$\mathrm{TPR}(t)$ $=\frac{|\mathrm{TP}(t)|}{|\mathrm{TP}(t)|+|\mathrm{FN}(t)|}=\frac{|\mathrm{TP}(t)|}{|\mathrm{P}|}$
@@
Appendix J Benchmarked Methods
@@
J.1 Spectral Normalized Gaussian Process
Spectral normalized Gaussian processes (SNGP) (Liu et al., 2020) use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty. For an input $x\in\mathcal{X}$ and number of classes $C$, they predict a $C$-variate Gaussian distribution
$\mathcal{N}\left(\bm{B}\bm{\phi}(x),\bm{\phi}(x)^{\top}\left(\bm{\Psi}^{\top}\bm{\Psi}+\bm{I}\right)^{-1}\bm{\phi}(x)\bm{I}_{C}\right),$ (93)
@@
  • $\bm{\Psi}^{\top}\bm{\Psi}$ is the (unnormalised) empirical covariance matrix of the pre-logits of the training set. This is calculated via accumulating the mini-batch estimates during the last epoch. [^2] As we use a cosine learning rate decay in all experiments, the model makes negligible changes in its pre-logit feature space in the last epoch. Thus, the empirical covariance matrix is approximately consistent.
The method applies spectral normalization to the hidden weights in each layer using a power iteration scheme with a single iteration per batch to obtain the largest singular value. Liu et al. (2020) claim this helps with input distance awareness.
J.2 Heteroscedastic Classifier
Heteroscedastic classifiers (HET) (Collier et al., 2021) construct a Gaussian distribution in the logit space to model per-input uncertainties:
@@
To learn the per-input covariance matrices from the training set, one has to construct a predictive estimate from $\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x))$ using any of the methods in Appendix F. This predictive estimate is then trained using a standard cross-entropy (NLL) loss.
@@
In Section 2.3, we show that when our constraint is satisfied, the off-diagonal terms of the covariance matrix do not matter for the predictive. This means that, in our framework, one can discard the low-rank term $\bm{V}(x)$ and only model the diagonal term $\bm{d}(x)$ without a decrease in expressivity. To keep comparisons fair and use the same backbone with the same number of parameters, we also only model $\bm{d}(x)$ for softmax-based predictives.
J.3 Laplace Approximation
The Laplace approximation (Daxberger et al., 2021) approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian. In its simplest form, it uses the maximum a posteriori (MAP) weights $\bm{\theta}_{\text{MAP}}\in\mathbb{R}^{P}$ as the mean and the inverse Hessian of the regularized loss over the training set ~$\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})=\mathcal{L}(\bm{\theta};\mathcal{D})+\lambda\|\bm{\theta}\|_{2}^{2}$ evaluated at the MAP as the covariance matrix:
@@
This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see (Tatzel et al., 2025).
For deep neural networks, the Hessian matrix is often replaced with the Generalized Gauss-Newton (GGN) matrix. The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature (Martens & Grosse, 2015) or low-rank approximations.
@@
$\mathcal{N}\left(\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)\bm{\theta}_{\text{MAP}},\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)\left(\bm{G}+\lambda\bm{I}_{P}\right)^{-1}\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)^{\top}\right),$ (97)
@@
We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates (Tatzel et al., 2025).
Appendix K CIFAR-10 Experiments
This appendix section repeats the experiments presented in the main paper on the CIFAR-10 dataset. For a detailed description of the experimental setup, refer to Appendix H. Appendix I describes the used tasks and metrics.
@@
  • What are the effects of changing the learning objective? (Section K.2)
  • Do we have to sacrifice performance for sample-free predictives? (Section K.1)
K.1 Quality of Sample-Free Predictives
Method Mean Std
@@
MC 100 0.0096 0.0013
MC 1000 0.0102 0.0016
MC 10 0.0120 0.0013
Mean Field 0.0121 0.0029
Laplace Bridge Predictive 0.5933 0.0105
@@
Similarly to the main paper, in this section, we investigate our first research question: whether there is a price to pay for sample-free predictives. Table 2 showcases the two best-performing (activation, method) pairs on the ECE metric and the CIFAR-10 dataset: Softmax and NormCDF Laplace. Mean Field (MF) is a strong alternative for sample-free predictives, but it has no guarantees and can fall behind MC sampling (see also Figure 1). When the constraint is satisfied, the analytic NormCDF predictive is exact. Empirically, our analytic predictives always perform on par with MC sampling.
K.2 Effects of Changing the Learning Objective
@@
As in the main paper, in this section, we use the best-performing predictive and estimator (see Appendix F) for softmax models and employ our methods with the analytic predictives.
K.2.1 Calibration and Proper Scoring
@@
We first evaluate calibration using the log probability scoring rule (Gneiting & Raftery, 2007) and the Expected Calibration Error (ECE) metric (Naeini et al., 2015). Figure 5 shows that on CIFAR-10, the score of our analytic predictives with implicit constraints (Sigmoid, NormCDF) are consistently better than the corresponding softmax results for all methods.
@@
Figure 11: CIFAR-10 ECE results. Analytic predictives with implicit ( ) or explicit ( ) constraints outperform Softmax ( ) on HET and SNGP. Laplace tunes its hyperparameters based on the ECE metric – NormCDF Laplace is the overall best method. Note the restricted $y$-limits for readability.
@@
Figure 11 shows that on CIFAR-10, our analytic predictives have a clear advantage on HET and SNGP. Laplace is a post-hoc method that tunes its hyperparameters on the ECE metric, hence its enhanced performance. Our NormCDF predictive is on par with Softmax.
K.2.2 Correctness Prediction
@@
Figure 12: CIFAR-10 log probability proper scoring results for the binary correctness prediction task. Analytic predictives with implicit ( ) constraints consistently outperform Softmax ( ) on all methods.
@@
Figure 12 shows that on the correctness prediction task, our analytic predictives with implicit constraints outperform all Softmax predictives across all methods, as measured by the log probability proper scoring rule.
K.2.3 Accuracy
@@
Figure 13: Analytic predictives do not sacrifice accuracy. CIFAR-10 accuracies. Analytic predictives with implicit ( ) or explicit ( ) constraints either outperform or are on par with Softmax ( ) across all methods.
@@
Analytic predictives do not sacrifice accuracy. Figure 13 evidences this claim on CIFAR-10: our analytic predictives either outperform or are on par with Softmax predictives. The most accurate method is Sigmoid HET. These results support the findings of Wightman et al. (2021) that showcase desirable training dynamics of the class-wise cross-entropy loss.
K.2.4 Out-of-Distribution Detection
@@
Figure 14: CIFAR-10C OOD detection AUROC results for severity level one. Across all methods, the best-performing predictive is analytic with an implicit constraint ( ).
@@
Finally, we consider the OOD detection task on a balanced mixture of ID (CIFAR-10) and OOD inputs. As OOD inputs, we consider corrupted CIFAR-10C samples. We use the AUROC metric to evaluate the methods’ performance. As shown in Figure 14, the best-performing method is Sigmoid SNGP, an analytic method. Generally, Softmax performs on par with our analytic predictives. Intuitively, separating ID and OOD samples does not require a fine-grained representation of uncertainty, unlike the ECE or proper scoring rules. Nevertheless, the analytic predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5).
Appendix L ImageNet Dirichlet Evaluation
This appendix section evaluates our moment-matched Dirichlet distributions on the OOD detection task (Appendix I). Table 3 shows the per-estimator performances of the three best-performing (method, activation) pairs on ImageNet: Sigmoid Laplace, NormCDF HET, and NormCDF Laplace. In all three cases, the best estimator is derived from the moment-matched Dirichlet distributions.
Table 3: Comparison of OOD AUROC results for different uncertainty estimators on the three best-performing (method, activation) pairs: Sigmoid Laplace, NormCDF HET, and NormCDF Laplace. The best estimator is derived from the moment-matched Dirichlet distributions in all three cases.
Method Mean Std
Sigmoid Laplace
Dirichlet Mutual Information 0.6388 0.0005
Analytic Predictive Entropy 0.6353 0.0003
MC 1000 Predictive Entropy 0.6323 0.0001
MC 100 Predictive Entropy 0.6323 0.0001
MC 1000 Expected Entropy 0.6321 0.0001
NormCDF HET
Dirichlet Expected Entropy 0.6353 0.0021
MC 1000 Expected Entropy 0.6281 0.0020
MC 100 Expected Entropy 0.6281 0.0020
MC 10 Expected Entropy 0.6281 0.0020
Analytic Predictive Entropy 0.6277 0.0021
NormCDF Laplace
Dirichlet Expected Entropy 0.6321 0.0009
MC 100 Predictive Entropy 0.6296 0.0012
MC 1000 Predictive Entropy 0.6296 0.0012
Closed-Form Predictive Entropy 0.6296 0.0012
MC 10 Predictive Entropy 0.6294 0.0012
Appendix M Further Out-of-Distribution Detection Results
@@
Figure 15 shows OOD detection results across all ImageNet-C severity levels. Our analytic predictives with implicit constraints consistently outperform Softmax.
(a) OOD detection AUROC with severity level one.
(b) OOD detection AUROC with severity level two.
(c) OOD detection AUROC with severity level three.
(d) OOD detection AUROC with severity level four.
(e) OOD detection AUROC with severity level five.
@@
Figure 15: The OOD detection performance of all methods increases steadily as we increase the severity of the perturbed half of the mixed dataset on the ImageNet validation dataset. Our analytic predictives consistently outperform Softmax.

### v2

@@ Appendix H Benchmark Metrics
H.1 Log Probability Proper Scoring Rule for the Predictive
First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on [26].
@@ H.1 Log Probability Proper Scoring Rule for the Predictive
H.2 Expected Calibration Error
To set up the required quantities for the Expected Calibration Error (ECE) metric [29], we follow the steps below, based on [26].
@@ H.2 Expected Calibration Error
The Expected Calibration Error (ECE) metric [29] is then defined as
@@
$\mathrm{conf}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{n\in B_{m}}\max_{1\leq c\leq C}f_{c}(x_{n}).$ (91)
@@
H.3 Binary Log Probability Proper Scoring Rule for Correctness Prediction
@@
$\ell\equiv\ell(x,y)=1\left(\max_{1\leq c\leq C}h_{c}(x)=y\right).$ (92)
@@
One can show that this is indeed a strictly proper scoring rule [26].
H.4 Accuracy
@@
H.5 Area Under the Receiver Operating Characteristic Curve for Out-of-Distribution Detection
@@
$\mathrm{TPR}(t)$ $=\frac{|\mathrm{TP}(t)|}{|\mathrm{TP}(t)|+|\mathrm{FN}(t)|}=\frac{|\mathrm{TP}(t)|}{|\mathrm{P}|}$ (96)
@@
Appendix I Benchmarked Methods
@@
I.1 Spectral Normalized Gaussian Process
Spectral normalized Gaussian processes (SNGP) [20] use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty. For an input $x\in\mathcal{X}$ and number of classes $C$, they predict a $C$-variate Gaussian distribution
$\mathcal{N}\left(\bm{B}\bm{\phi}(x),\bm{\phi}(x)^{\top}\left(\bm{\Psi}^{\top}\bm{\Psi}+\bm{I}\right)^{-1}\bm{\phi}(x)\bm{I}_{C}\right)$ (97)
@@ I.1 Spectral Normalized Gaussian Process
  • $\bm{\Psi}^{\top}\bm{\Psi}$ is the (unnormalised) empirical covariance matrix of the pre-logits of the training set. This is calculated by accumulating the mini-batch estimates during the last epoch. [^3] As we use a cosine learning rate decay in all experiments, the model makes negligible changes in its pre-logit feature space in the last epoch. Thus, the empirical covariance matrix is approximately consistent.
The method applies spectral normalization to the hidden weights in each layer using a power iteration scheme with a single iteration per batch to obtain the largest singular value. Liu et al. [20] claim this helps with input distance awareness.
I.2 Heteroscedastic Classifier
Heteroscedastic classifiers (HET) [5] construct a Gaussian distribution in the logit space to model per-input uncertainties:
@@ I.2 Heteroscedastic Classifier
To learn the per-input covariance matrices from the training set, one has to construct a predictive estimate from $\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x))$ using any of the methods in Appendix E. This predictive estimate is then trained using a standard cross-entropy (NLL) loss.
@@
The off-diagonal terms of the covariance matrix do not affect the approximate predictive (Eq. 12). This means that, in our framework, one can discard the low-rank term $\bm{V}(x)$ and only model the diagonal term $\bm{d}(x)$ without a decrease in expressivity. To keep comparisons fair and use the same backbone with the same number of parameters, we also only model $\bm{d}(x)$ for softmax-based predictives.
I.3 Laplace Approximation
The Laplace approximation [8] approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian. In its simplest form, it uses the maximum a posteriori (MAP) weights $\bm{\theta}_{\text{MAP}}\in\mathbb{R}^{P}$ as the mean and the inverse Hessian of the regularised loss over the training set ~$\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})=\mathcal{L}(\bm{\theta};\mathcal{D})+\lambda\|\bm{\theta}\|_{2}^{2}$ evaluated at the MAP as the covariance matrix:
@@ I.3 Laplace Approximation
This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see [35].
For deep neural networks, the Hessian matrix is often replaced with the Generalized Gauss-Newton (GGN) matrix. The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature [24] or low-rank approximations.
@@
$\mathcal{N}\left(\bm{f}(x,\bm{\theta}_{\text{MAP}}),\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)\left(\bm{G}+\lambda\bm{I}_{P}\right)^{-1}\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)^{\top}\right),$ (101)
@@
We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates [35].
Appendix J CIFAR-10 Experiments
This appendix section repeats the experiments presented in the main paper on the CIFAR-10 dataset. For a detailed description of the experimental setup, refer to Appendix G. Appendix H describes the used tasks and metrics.
@@
  • What are the effects of changing the learning objective? (Section J.2)
  • Do we have to sacrifice performance for sample-free predictives? (Section J.1)
J.1 Quality of Sample-Free Predictives
Table J.1: Comparison of ECE results for different predictives using a fixed Laplace backbone.
Method Mean $\pm$ 2 std
@@
MC 100 0.0096 $\pm$ 0.0013
MC 1000 0.0102 $\pm$ 0.0016
MC 10 0.0120 $\pm$ 0.0013
Mean Field 0.0121 $\pm$ 0.0029
Laplace Bridge Predictive 0.5933 $\pm$ 0.0105
@@
Similarly to the main paper, in this section, we investigate our first research question: whether there is a price to pay for sample-free predictives. Table J.1 showcases the two best-performing (activation, method) pairs on the ECE metric and the CIFAR-10 dataset: Softmax and NormCDF Laplace. Mean Field (MF) is a strong alternative for sample-free predictives, but it has no guarantees and can fall behind MC sampling (see also Fig. 1). Empirically, our closed-form predictives always perform on par with MC sampling.
J.2 Effects of Changing the Learning Objective
@@
As in the main paper, in this section, we use the best-performing predictive and estimator (see Appendix E) for softmax models and employ our methods with the closed-form predictives.
J.2.1 Calibration and Proper Scoring
@@
We first evaluate calibration using the log probability scoring rule [12] and the Expected Calibration Error (ECE) metric [29]. Fig. 3(b) shows that on CIFAR-10, the score of our closed-form predictives (Sigmoid, NormCDF) are consistently better than the corresponding softmax results for all methods.
@@
Figure J.1: CIFAR-10 ECE results. Our closed-form predictives ( ) outperform Softmax ( ) on HET and SNGP. Laplace tunes its hyperparameters based on the ECE metric – NormCDF Laplace is the overall best method. Note the restricted $y$-limits for readability.
@@
Fig. J.1 shows that on CIFAR-10, our closed-form predictives have a clear advantage on HET and SNGP. Laplace is a post-hoc method that tunes its hyperparameters on the ECE metric, hence its enhanced performance. Our NormCDF predictive is on par with Softmax.
J.2.2 Correctness Prediction
@@
Figure J.2: CIFAR-10 log probability proper scoring results for the binary correctness prediction task. Our closed-form predictives ( ) consistently outperform Softmax ( ) on all methods.
@@
Fig. J.2 shows that on the correctness prediction task, our closed-form predictives outperform all Softmax predictives across all methods, as measured by the log probability proper scoring rule.
J.2.3 Accuracy
@@
Figure J.3: Closed-form predictives do not sacrifice accuracy. CIFAR-10 accuracies. Our closed-form predictives ( ) either outperform or are on par with Softmax ( ) across all methods.
@@
Closed-form predictives do not sacrifice accuracy. Fig. J.3 evidences this claim on CIFAR-10: our closed-form predictives either outperform or are on par with Softmax predictives. The most accurate method is Sigmoid HET. These results support the findings of Wightman et al. [38] that showcase desirable training dynamics of the class-wise cross-entropy loss.
J.2.4 Out-of-Distribution Detection
@@
Figure J.4: CIFAR-10C OOD detection AUROC results for severity level one. Across all methods, the best-performing predictive is closed-form ( ).
@@
Finally, we consider the OOD detection task on a balanced mixture of ID (CIFAR-10) and OOD inputs. As OOD inputs, we consider corrupted CIFAR-10C samples. We use the AUROC metric to evaluate the methods’ performance. As shown in Fig. J.4, the best-performing method is Sigmoid SNGP, a closed-form method. Generally, Softmax performs on par with our closed-form predictives. Intuitively, separating ID and OOD samples does not require a fine-grained representation of uncertainty, unlike the ECE or proper scoring rules. Nevertheless, the closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5).
Appendix K Additional Results
K.1 Closed-form Softmax Predictives
One can directly apply Eq. 10 to a neural network trained with the softmax cross-entropy loss in Eq. 4 to obtain closed-form predictives. However, we empirically found this approach to decrease performance, as the denominator of Eq. 10 increased to the order of billions during training. Considering a multivariate normal random variable $\bm{x}\sim\mathcal{N}(\bm{\mu},\Sigma)$ representing a logit distribution,
$\mathrm{Var}\left(\sum_{i=1}^{C}e^{x_{i}}\right)\geq\mathrm{Var}(e^{x_{j}})=e^{2\mu_{j}+\Sigma_{jj}}\underbrace{(e^{\Sigma_{jj}}-1)}_{>0},$ (102)
i.e., when $e^{\mu_{j}}$ is large for some $j\in\{1,\dots,C\}$, the variance of the denominator necessarily explodes, violating our assumption.
To mitigate this issue, one may optimize the regularised cross-entropy loss
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})=\tilde{\mathcal{L}}((x_{n},c_{n})_{n=1}^{N})+\lambda\sum_{n=1}^{N}\left(\sum_{c=1}^{C}\exp(f_{c}(x_{n}))-1\right)^{2}$ (103)
or the more numerically stable version
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})=\tilde{\mathcal{L}}((x_{n},c_{n})_{n=1}^{N})+\lambda\sum_{n=1}^{N}\left(\log\sum_{c=1}^{C}\exp(f_{c}(x_{n}))\right)^{2}$ (104)
for an appropriately tuned $\lambda$ hyperparameter. The latter formulation can be stably trained even at an ImageNet scale. However, there are fundamental limitations to the predictive given by Eq. 10. Namely, for isotropic logit-space Gaussian distributions, $\nicefrac{{\sigma^{2}(x)}}{{2}}$ introduces a constant shift in the logits, under which the softmax activation function is invariant. Therefore, the predictive approximation collapses into the MAP prediction. The SNGP method predicts such isotropic logit-space Gaussians. Further, we empirically found that the Laplace method’s predictives for models trained with the regularised cross-entropy loss were also approximately isotropic; thus, this integral approximation yielded diminishing returns.
For completeness, we share the ImageNet and CIFAR-10 results of $\varphi=\exp$ in Tables K.1 and K.2.
Table K.1: Performance metrics of $\varphi=\exp$ on the CIFAR-10 validation dataset. We report the mean and two standard deviations.
Metric Exp SNGP Exp HET Exp Laplace
NLL $\phantom{-}0.376\pm 0.007$ $\phantom{-}0.353\pm 0.059$ $\phantom{-}0.370\pm 0.024$
ECE $\phantom{-}0.041\pm 0.006$ $\phantom{-}0.029\pm 0.007$ $\phantom{-}0.045\pm 0.007$
Log Prob. $-0.255\pm 0.005$ $-0.238\pm 0.038$ $-0.255\pm 0.016$
Accuracy $\phantom{-}0.884\pm 0.013$ $\phantom{-}0.888\pm 0.023$ $\phantom{-}0.893\pm 0.007$
AUROC $\phantom{-}0.592\pm 0.006$ $\phantom{-}0.603\pm 0.009$ $\phantom{-}0.596\pm 0.011$
Table K.2: Performance metrics of $\varphi=\exp$ on the ImageNet validation dataset. We report the mean and two standard deviations.
Metric Exp SNGP Exp HET Exp Laplace
NLL $\phantom{-}0.866\pm 0.031$ $\phantom{-}0.929\pm 0.042$ $\phantom{-}0.986\pm 0.149$
ECE $\phantom{-}0.058\pm 0.004$ $\phantom{-}0.022\pm 0.007$ $\phantom{-}0.058\pm 0.013$
Log Prob. $-0.380\pm 0.009$ $-0.403\pm 0.011$ $-0.402\pm 0.014$
Accuracy $\phantom{-}0.784\pm 0.001$ $\phantom{-}0.779\pm 0.001$ $\phantom{-}0.785\pm 0.002$
AUROC $\phantom{-}0.606\pm 0.001$ $\phantom{-}0.611\pm 0.001$ $\phantom{-}0.615\pm 0.003$
K.2 Vision Transformer Results on ImageNet
Table K.3 shows results on ViT Little [11] backbones from the timm [37] library on the ImageNet validation set. As in the main paper, the hyperparameters are optimized using a ten-step Bayesian hyperparameter optimization scheme of Weights and Biases.
Table K.3: NLL and accuracy metrics of ViT Little models on the ImageNet validation set. Error bars represent two standard deviations.
Method NLL Accuracy (%)
Laplace
Softmax $0.81988\pm 0.0020$ $79.598\pm 0.145$
NormCDF $0.79603\pm 0.0087$ $80.678\pm 0.042$
Sigmoid $0.79479\pm 0.0002$ $80.104\pm 0.206$
HET
Softmax $0.87096\pm 0.0242$ $78.604\pm 0.378$
NormCDF $0.79106\pm 0.0075$ $80.514\pm 0.425$
Sigmoid $0.88175\pm 0.0356$ $78.738\pm 0.324$
GP
Softmax $0.86724\pm 0.0229$ $78.898\pm 0.187$
NormCDF $0.76974\pm 0.0088$ $80.694\pm 0.575$
Sigmoid $0.81178\pm 0.0115$ $79.980\pm 0.412$
K.3 CIFAR-100 Results
We repeat the NLL and accuracy evaluation on CIFAR-100 using WideResNet-28-5 models following a ten-step Bayesian hyperparameter sweep on Weights and Biases. Results are shown in Table K.4.
Table K.4: NLL and accuracy metrics of WideResNet-28-5 models on the CIFAR-100 test set. Error bars represent two standard deviations.
Method NLL Accuracy (%)
Laplace
Softmax $0.92176\pm 0.0128$ $78.225\pm 0.412$
NormCDF $0.88823\pm 0.0094$ $78.900\pm 0.325$
Sigmoid $0.94209\pm 0.0136$ $78.050\pm 0.436$
HET
Softmax HET $0.98513\pm 0.0254$ $77.050\pm 0.228$
NormCDF HET $0.94643\pm 0.0022$ $78.117\pm 0.286$
Sigmoid HET $0.95251\pm 0.0142$ $77.583\pm 0.175$
SNGP
Softmax SNGP $0.97156\pm 0.0348$ $77.750\pm 0.462$
NormCDF SNGP $0.90011\pm 0.0106$ $79.010\pm 0.298$
Sigmoid SNGP $0.99996\pm 0.0265$ $78.350\pm 0.215$
K.4 Alignment with the True Predictives
Table K.5 shows the divergence of predictives using various approximation techniques from the true predictive estimated using $10,000$ Monte Carlo samples. Our closed-form predictives, NormCDF and Sigmoid, are favorable to both the mean field and Laplace bridge approximations.
Table K.5: Kullback-Leibler (KL) divergence to the true predictive distributions (estimated using $10,000$ Monte Carlo samples) for different approximation methods on the ImageNet validation set using ViT Little backbones. The mean KL divergence is calculated over the validation set. Error bars represent two standard deviations over independently trained models using different seeds.
Method KL Divergence to True Predictive
NormCDF $0.0057\pm 0.0008$
Sigmoid $0.0064\pm 0.0009$
Softmax mean field $0.0330\pm 0.0042$
Softmax Laplace bridge $1.7900\pm 0.1254$
K.5 BCE Loss Performance Gains
Table K.6 shows that the Softmax and Sigmoid activation functions (and corresponding losses) already show improved performance on the vanilla models compared to softmax, highlighting that the performance gains we observe cannot only be attributed to our closed-form predictive approximations but also the favorable training dynamics of the class-wise BCE loss.
Table K.6: NLL results on ImageNet using ResNet-50 backbones and different activation functions. Error bars represent two standard deviations.
Activation Function NLL
NormCDF $0.9345\pm 0.0072$
Softmax $0.9369\pm 0.0025$
Sigmoid $0.9146\pm 0.0046$
K.6 Further Out-of-Distribution Detection Results
@@
Fig. K.1 shows OOD detection results across all ImageNet-C severity levels. Our closed-form predictives consistently outperform Softmax.
@@
Figure K.1: The OOD detection performance of all methods increases steadily as we increase the severity of the perturbed half of the mixed dataset on the ImageNet validation dataset. Our closed-form predictives consistently outperform Softmax.